# Application of Operator Theory for the Collatz Conjecture

### v1

@@ 4.3 Formulation by the Cuntz algebra
While we defined C*-algebras as subalgebras of B(H) for some Hilbert space H in Section 2, they can be characterized abstractly without using Hilbert spaces (see e.g., [3], [4] for more details). The notions of projections, isometries, and partial isometries can be defined in the same way. The Cuntz algebras, which were introduced by Cuntz in [5], are typical examples of such abstract C*-algebras. They are defined as universal C*-algebras given by generators and relations as follows.
@@
Cuntz showed that $\mathcal{O}_{n}$ is simple as a $C^{*}$-algebra in [5], and so, for any isometries $S_{1},S_{2},\cdots,S_{n}\in\mathcal{B}(\mathcal{H})$ satisfying the relation above, $C^{*}(S_{1},S_{2},\cdots,S_{n})$ is isomorphic to $\mathcal{O}_{n}$.
@@
$N_{1}$ =$=\{n\in\mathbb{N}\mid n\equiv 1,5\pmod{6}\},$ (35)
$N_{2}$ $=f(N_{1})=\{n\in\mathbb{N}\mid n\equiv 4,16\pmod{18}\}.$ (36)
@@
$T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$ (38)
$T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2}.\end{cases}$ (39)
@@
$\overline{C^{*}(S_{1},S_{2})e_{n}}=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2},m\sim n\}}.$ (40)
@@
It suffices to show that $\overline{C^{*}(T_{1},T_{2})e_{n}}=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2},m\sim n\}}$. The way to prove is the same as the proof of Lemma 4.2.2. ∎
@@
$\overline{C^{*}(S_{1},S_{2})e_{1}}$ $=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2},m\sim 1\}}=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2}\}}=\mathcal{H}.$
@@
00 0 $\neq\left\langle Qe_{1},e_{n}\right\rangle=\left\langle QT_{1}^{*}T_{1}e_{1},e_{n}\right\rangle=\left\langle QT_{1}e_{1},T_{1}e_{n}\right\rangle$
$=\left\langle Qe_{4},T_{1}e_{n}\right\rangle=\left\langle QT_{2}^{*}T_{2}e_{4},T_{1}e_{n}\right\rangle=\left\langle QT_{2}e_{4},T_{2}T_{1}e_{n}\right\rangle=\left\langle Qe_{1},T_{2}T_{1}e_{n}\right\rangle.$
@@
$P^{2k}(n)\in N_{1}.$ (42)
If the Collatz conjecture holds, then $1\in\operatorname{orb}(n;P)$ by Proposition 4.3.3. It follows that $P^{2k}(n)\in\{1,4\}$ for sufficiently large $k\in\mathbb{N}$. Since $P^{2k}(n)\in N_{1}$, $P_{2k}(n)=1$. Then,
$(T_{2}T_{1})^{k}e_{n}=e_{1}.$ (43)
@@
But, this contradicts the assumption $n\in(N_{1}\cup N_{2})\backslash\{1\}$.
Thus, $\left\langle Qe_{1},e_{n}\right\rangle=0$ for every $n\in(N_{1}\cup N_{2})\backslash\{1\}$. Then, it holds that $Qe_{1}=\left\langle Qe_{1},e_{1}\right\rangle e_{1}=\left\lVert Qe_{1}\right\rVert^{2}e_{1}$, and hence $Qe_{1}\in\{0,e_{1}\}$. ∎
@@ 5.1 Bounded condition
When there exist disjoint subsets $X_{1},X_{2},\cdots,X_{k}$ of $X$, where $k\in\mathbb{N}$ such that:
@@
    $f_{q}(n)=\begin{cases}qn+1,&n:\text{odd},\\ n/2,&n:\text{even}.\end{cases}$ (44)
@@
    $f_{d}(n)=\begin{cases}3n+d,&n:\text{odd},\\ n/2,&n:\text{even}.\end{cases}$ (45)
@@
$T\left(\sum_{x\in F}a_{x}e_{x}\right)=\sum_{x\in F}a_{x}Te_{x},\text{ where $F$ is a finite subset of $X$, }a_{x}\in\mathbb{C}.$ (47)
According to the bounded condition for $f$, $f^{-1}(x)$ includes at most $k$ elements for any $x\in X$ and so $\left\lvert\sum_{y\in f^{-1}(x)}a_{y}\right\rvert^{2}\leq(\sum_{y\in f^{-1}(x)}\left\lvert a_{y}\right\rvert)^{2}\leq k\sum_{y\in f^{-1}(x)}\left\lvert a_{y}\right\rvert^{2}$. Since
$\sum_{x\in X}\left\lvert\sum_{y\in f^{-1}(x)}a_{y}\right\rvert^{2}\leq\sum_{x\in X}k\sum_{y\in f^{-1}(x)}\left\lvert a_{y}\right\rvert^{2}=k\sum_{x\in X}\left\lvert a_{x}\right\rvert^{2}=k\left\lVert x\right\rVert^{2},$
there exists $\sum_{x\in X}\left(\sum_{y\in f^{-1}(x)}a_{x}\right)e_{x}\in\mathcal{H}$. Thus, we can define $Tx$ as
$Tx=\sum_{x\in X}a_{x}Te_{x}=\sum_{x\in X}a_{x}e_{f(x)}=\sum_{x\in X}\left(\sum_{y\in f^{-1}(x)}a_{y}\right)e_{x},$ (48)
@@
$\overline{C^{*}(T)e_{x}}\subseteq\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}.$ (49)
@@
$T^{*}e_{z}=\sum_{w\in X}\left\langle T^{*}e_{z},e_{w}\right\rangle e_{w}=\sum_{w\in X}\left\langle e_{z},Te_{w}\right\rangle e_{w}=\sum_{w\in X}\left\langle e_{z},e_{f(w)}\right\rangle e_{w}=\sum_{w\in f^{-1}(z)}e_{w}.$
@@
When $e_{x}$ is a cyclic vector for $C^{*}(T)$, $\mathcal{H}=\overline{C^{*}(T)e_{x}}\subseteq\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}$ by Lemma 5.1.4. It follows that $y\sim x$ for every $y\in X$. Pick any $y,z\in X$. Since $y\sim x$ and $z\sim x$, it holds that $y\sim z$. ∎
@@
$T_{i}e_{x}=\begin{cases}e_{f(x)},&x\in X_{i},\\ 0,&x\notin X_{i},\end{cases}$ (50)
@@
$\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})e_{x}}=\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}.$ (51)
@@
$e_{y}$ $=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{m}})^{*}e_{f^{m}(y)}=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{m}})^{*}e_{f^{n}(x)}=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{m}})^{*}(T_{j_{1}}T_{j_{2}}\cdots T_{j_{n}})e_{x}.$
@@ 5.2 Separating condition
$maps-to11\mapsto 2\mapsto 3\mapsto\cdots\mapsto n-1\mapsto n\$ 1 . (55)
@@
$(i_{\sigma^{j}(1)},i_{\sigma^{j}(2)},\cdots,i_{\sigma^{j}(n)})\neq(i_{1},i_{2},\cdots,i_{n})$ (56)
@@
    For the minimum $n$ as above, the $n$-tuple $(i_{1},i_{2},\cdots,i_{n})\in\{1,2,,\cdots,k\}^{*}$, such that $f^{j-1}(x)\in X_{i_{j}}$ where $1\leq j\leq n$, is aperiodic,
@@
$T_{j_{m}}T_{j_{m-1}}\cdots T_{j_{1}}e_{y}=e_{x}.$ (59)
Let $b=T_{j_{m}}T_{j_{m-1}}\cdots T_{j_{1}}a$. It follows that $\left\langle b,e_{x}\right\rangle\neq 0$ and $b\in\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})a}$. Since $\left\lVert T_{I}\right\rVert\leq 1$,
$\left\lVert T_{I}^{m}b\right\rVert\geq\left\lVert T_{I}^{m+1}b\right\rVert$ (60)
@@
$\mathcal{M}=\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})e_{x}}\subseteq\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})a}\subseteq\mathcal{M}.$ (61)
@@
$\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})e_{x}}=\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}=\overline{\operatorname{span}\{e_{y}\mid y\in X\}}=\mathcal{H}.$ (62)
@@
$\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})a}=\mathcal{H}$ (63)
for all $\0a\in\mathcal{H}\backslash\{0\}$. Then every $\0a\in\mathcal{H}\backslash\{0\}$ is a cyclic vector for $C^{*}(T_{1},T_{2},\cdots,T_{k})$ and this implies that $C^{*}(T_{1},T_{2},\cdots,T_{k})$ has no non-trivial reducing subspaces. ∎
@@ 5.3 Cuntz-Krieger condition
$A(j,i)=\begin{cases}1,&i\in J,\\ 0,&o.w.\end{cases}$ (64)
@@ 5.4 Transformations of dynamical systems
    $T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$ (65)
    $T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2},\end{cases}$ (66)
    and $S_{1},S_{2}\in\mathcal{B}(\mathcal{H})$ as $S_{1}=T_{1}^{*}T_{2}^{*},S_{2}=T_{2}^{*}$. Then, $C^{*}(T_{1},T_{2})=C^{*}(S_{1},S_{2})$ is isomorphic to the Cuntz algebra $\mathcal{O}_{2}$.
@@
    $N_{2}=f_{q}(N_{1})\subseteq\{n\in\mathbb{N}\mid n\equiv 1+q,1+3q,1+5q,\cdots,1+(2q-1)q\pmod{2q^{2}}\}.$
@@
    $T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$ (67)
    $T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2},\end{cases}$ (68)
    and $S_{1},S_{2}\in\mathcal{B}(\mathcal{H})$ as $S_{1}=T_{1}^{*}T_{2}^{*},S_{2}=T_{2}^{*}$. Then, $C^{*}(T_{1},T_{2})=C^{*}(S_{1},S_{2})$ is isomorphic to the Cuntz algebra $\mathcal{O}_{2}$.
@@
    $T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$ (69)
    $T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2},\end{cases}$ (70)
    and $S_{1},S_{2}\in\mathcal{B}(\mathcal{H})$ as $S_{1}=T_{1}^{*}T_{2}^{*},S_{2}=T_{2}^{*}$. Then, $C^{*}(T_{1},T_{2})=C^{*}(S_{1},S_{2})$ is isomorphic to the Cuntz algebra $\mathcal{O}_{2}$.

### v2

@@ 4.3 Formulation by the Cuntz algebra
While we defined C*-algebras as subalgebras of $\mathcal{B}(\mathcal{H})$ for some Hilbert space H in Section 2, they can be characterized abstractly without using Hilbert spaces (see e.g., [4], [3] for more details). The notions of projections, isometries, and partial isometries can be defined in the same way. The Cuntz algebras, which were introduced by Cuntz in [5], are typical examples of such abstract C*-algebras. They are defined as universal C*-algebras given by generators and relations as follows.
@@
Cuntz showed that $\mathcal{O}_{n}$ is simple as a $C^{*}$-algebra in [5], and so, for any isometries $S_{1},S_{2},\cdots,S_{n}\in\mathcal{B}(\mathcal{H})$ satisfying the relation above, it holds that $C^{*}(S_{1},S_{2},\cdots,S_{n})$ is isomorphic to $\mathcal{O}_{n}$.
@@
$N_{1}$ =$=\{n\in\mathbb{N}\mid n\equiv 1,5\pmod{6}\},$
$N_{2}$ $=f(N_{1})=\{n\in\mathbb{N}\mid n\equiv 4,16\pmod{18}\}.$
@@
$T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$
$T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2}.\end{cases}$
@@
$\overline{C^{*}(S_{1},S_{2})e_{n}}=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2},m\sim n\}}.$
@@
It suffices to show that
$\overline{C^{*}(T_{1},T_{2})e_{n}}=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2},m\sim n\}}.$
The way to prove is the same as the proof of Lemma 4.2.2. ∎
@@
$\overline{C^{*}(S_{1},S_{2})e_{1}}$ $=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2},m\sim 1\}}$
$=\overline{\operatorname{span}\{e_{m}\mid m\in N_{1}\cup N_{2}\}}=\mathcal{H}.$
@@
We remark that $T_{2}T_{1}e_{1}=e_{1}$ and $e_{1}=T_{1}^{*}T_{2}^{*}e_{1}$.
Let $Q\in C^{*}(S_{1},S_{2})^{\prime}=C^{*}(T_{1},T_{2})^{\prime}$ be a projection and assume $Qe_{1}\neq 0$. We will show $Qe_{1}=e_{1}$.
@@
$0\neq\left\langle Qe_{1},e_{n}\right\rangle=\left\langle QT_{1}^{*}T_{2}^{*}e_{1},e_{n}\right\rangle=\left\langle T_{1}^{*}T_{2}^{*}Qe_{1},e_{n}\right\rangle=\left\langle Qe_{1},T_{2}T_{1}e_{n}\right\rangle.$
@@
$P^{2k}(n)\in N_{1}.$
If the Collatz conjecture holds, then $1\in\operatorname{orb}(n;P)$ by Proposition 4.3.3. It follows that $P^{2k}(n)\in\{1,4\}$ for sufficiently large $k\in\mathbb{N}$. Since $P^{2k}(n)\in N_{1}$, $P^{2k}(n)=1$. Then,
$(T_{2}T_{1})^{k}e_{n}=e_{1}.$
@@
But, this contradicts the assumption $n\in(N_{1}\cup N_{2})\backslash\{1\}$. Thus, $\left\langle Qe_{1},e_{n}\right\rangle=0$ for every $n\in(N_{1}\cup N_{2})\backslash\{1\}$.
Hence,
$Qe_{1}=\sum_{n\in N_{1}\cup N_{2}}\left\langle Qe_{1},e_{n}\right\rangle e_{n}=\left\langle Qe_{1},e_{1}\right\rangle e_{1}=\left\lVert Qe_{1}\right\rVert^{2}e_{1}.$
Since $Qe_{1}\neq 0$, it holds that $\left\lVert Qe_{1}\right\rVert=1$ and $Qe_{1}=e_{1}$. ∎
@@ 5.1 Bounded condition
When there exist disjoint subsets $X_{1},\\ X_{2},\cdots,X_{k}$ of $X$, where $k\in\mathbb{N}$ such that:
@@
    $f_{q}(n)=\begin{cases}qn+1,&n:\text{odd},\\ n/2,&n:\text{even}.\end{cases}$
@@
    $f_{d}(n)=\begin{cases}3n+d,&n:\text{odd},\\ n/2,&n:\text{even}.\end{cases}$
@@
$T\left(\sum_{x\in F}a_{x}e_{x}\right)=\sum_{x\in F}a_{x}Te_{x},\text{ where $F$ is a finite subset of $X$, }a_{x}\in\mathbb{C}.$
According to the bounded condition for $f$, $f^{-1}(x)$ includes at most $k$ elements for any $x\in X$ and so
$\left\lvert\sum_{y\in f^{-1}(x)}a_{y}\right\rvert^{2}\leq(\sum_{y\in f^{-1}(x)}\left\lvert a_{y}\right\rvert)^{2}\leq k\sum_{y\in f^{-1}(x)}\left\lvert a_{y}\right\rvert^{2}.$
Since
$\sum_{x\in X}\left\lvert\sum_{y\in f^{-1}(x)}a_{y}\right\rvert^{2}\leq\sum_{x\in X}k\sum_{y\in f^{-1}(x)}\left\lvert a_{y}\right\rvert^{2}=k\sum_{x\in X}\left\lvert a_{x}\right\rvert^{2},$
there exists $\sum_{x\in X}\left(\sum_{y\in f^{-1}(x)}a_{x}\right)e_{x}\in\mathcal{H}$ for every $a=\sum_{x\in X}a_{x}e_{x}\in\mathcal{H}$. Thus, we can define $Ta$ as
$Ta=\sum_{x\in X}a_{x}Te_{x}=\sum_{x\in X}a_{x}e_{f(x)}=\sum_{x\in X}\left(\sum_{y\in f^{-1}(x)}a_{y}\right)e_{x},$
@@
$\overline{C^{*}(T)e_{x}}\subseteq\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}.$
@@
$T^{*}e_{z}$ $=\sum_{w\in X}\left\langle T^{*}e_{z},e_{w}\right\rangle e_{w}=\sum_{w\in X}\left\langle e_{z},Te_{w}\right\rangle e_{w}$
$=\sum_{w\in X}\left\langle e_{z},e_{f(w)}\right\rangle e_{w}=\sum_{w\in f^{-1}(z)}e_{w}.$
@@
When $e_{x}$ is a cyclic vector for $C^{*}(T)$,
$\mathcal{H}=\overline{C^{*}(T)e_{x}}\subseteq\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}$
by Lemma 5.1.4. It follows that $y\sim x$ for every $y\in X$. Pick any $y,z\in X$. Since $y\sim x$ and $z\sim x$, it holds that $y\sim z$. ∎
@@
$T_{i}e_{x}=\begin{cases}e_{f(x)},&x\in X_{i},\\ 0,&x\notin X_{i},\end{cases}$
@@
$\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})e_{x}}=\overline{\operatorname{span}\{e_{y}\mid y\in X,y\sim x\}}.$
@@
$e_{y}$ $=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{m}})^{*}e_{f^{m}(y)}=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{m}})^{*}e_{f^{n}(x)}$
$=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{m}})^{*}(T_{j_{1}}T_{j_{2}}\cdots T_{j_{n}})e_{x}.$
@@ 5.2 Separating condition
$maps-to11\mapsto 2\mapsto 3\mapsto\cdots\mapsto n-1\mapsto n\$ 1 .
@@
$(i_{\sigma^{j}(1)},i_{\sigma^{j}(2)},\cdots,i_{\sigma^{j}(n)})\neq(i_{1},i_{2},\cdots,i_{n})$
@@
    For the minimum $n$ as above, the $n$-tuple $(i_{1},i_{2},\cdots,i_{n})\in\{1,2,\cdots,k\}^{*}$, such that $f^{j-1}(x)\in X_{i_{j}}$ where $1\leq j\leq n$, is aperiodic,
@@
$T_{j_{m}}T_{j_{m-1}}\cdots T_{j_{1}}e_{y}=e_{x}.$
Let $b=T_{j_{m}}T_{j_{m-1}}\cdots T_{j_{1}}a$. It follows that $\left\langle b,e_{x}\right\rangle\neq 0$, $b\in\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})a}$. Since $\left\lVert T_{I}\right\rVert\leq 1$,
$\left\lVert T_{I}^{m}b\right\rVert\geq\left\lVert T_{I}^{m+1}b\right\rVert$
@@
$\mathcal{M}=\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})e_{x}}\subseteq\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})a}\subseteq\mathcal{M}.$
@@
$\overline{C^{*}(T_{1},T_{2},\cdots,T_{k})a}=\mathcal{H}$
for every $\0a\in\mathcal{H}\backslash\{0\}$. Then $a$ is a cyclic vector for $C^{*}(T_{1},T_{2},\cdots,T_{k})$ and this implies that $C^{*}(T_{1},T_{2},\cdots,T_{k})$ has no non-trivial reducing subspaces. ∎
@@ 5.3 Cuntz-Krieger condition
$A(j,i)=\begin{cases}1,&i\in J,\\ 0,&o.w.\end{cases}$
@@ 5.4 Transformations of dynamical systems
    $T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$
    $T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2},\end{cases}$
    and $S_{1},S_{2}\in\mathcal{B}(\mathcal{H})$ as $S_{1}=T_{1}^{*}T_{2}^{*},S_{2}=T_{2}^{*}$. It follows that, $C^{*}(T_{1},T_{2})=C^{*}(S_{1},S_{2})$ is isomorphic to the Cuntz algebra $\mathcal{O}_{2}$.
@@
    $N_{2}=f_{q}(N_{1})\subseteq$
    $\{n\in\mathbb{N}\mid n\equiv 1+q,1+3q,1+5q,\cdots,1+(2q-1)q\pmod{2q^{2}}\}.$
@@
    $T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$
    $T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2},\end{cases}$
    and $S_{1},S_{2}\in\mathcal{B}(\mathcal{H})$ as $S_{1}=T_{1}^{*}T_{2}^{*},S_{2}=T_{2}^{*}$. It follows that, $C^{*}(T_{1},T_{2})=C^{*}(S_{1},S_{2})$ is isomorphic to the Cuntz algebra $\mathcal{O}_{2}$.
@@
    $T_{1}e_{n}$ =$=\begin{cases}e_{P(n)},&n\in N_{1},\\ 0,&n\in N_{2},\end{cases}$
    $T_{2}e_{n}$ =$=\begin{cases}0,&n\in N_{1},\\ e_{P(n)},&n\in N_{2},\end{cases}$
    and $S_{1},S_{2}\in\mathcal{B}(\mathcal{H})$ as $S_{1}=T_{1}^{*}T_{2}^{*},S_{2}=T_{2}^{*}$. It follows that, $C^{*}(T_{1},T_{2})=C^{*}(S_{1},S_{2})$ is isomorphic to the Cuntz algebra $\mathcal{O}_{2}$.